a 250-N crate is pulled along a horizontal sidewalk at a constant speed by a rope which makes an angle of 40° with the sidewalk. If a force of 150 N is applied on the rope, what is the coefficient of the sliding friction?

1
Expert's answer
2021-11-16T09:52:42-0500

Write the balance of forces for motion at constant speed:


"-\mu(W-F\sin\theta)+F\cos\theta=0,\\\space\\
\mu=\frac{F\cos\theta}{W-F\sin\theta}=0.75."